A rectangle is 10 cm longer than it is wide. A line segment cuts the area enclosed into two pieces, one of which is a square. The area of the rectangle is 118cm squared more than the area of the square. What is the width of the rectangle?
Answer by ankor@dixie-net.com(12693)

A rectangle is 10 cm longer than it is wide. A line segment cuts the area enclosed into two pieces, one of which is a square. The area of the rectangle is 118cm squared more than the area of the square. What is the width of the rectangle?
:
Let x = the width of the rectangle; (x+10) = the length
:
The square sides would = x, it's area would be x^2
:
"The area of the rectangle is 118 cm squared more than the area of the square."
Means the Area of the rectangle = (x^2 + 118)
:
We also know that the Area of the rectangle = x(x+10)= (x^2 + 10x)
:
Therefore:
x^2 + 10x = x^2 + 118
x^2 - x^2 + 10x = 118
10x = 118
x = 11.8 cm is the width
:
:
Check:
11.8^2 + 118 = 257.24
11.8 * 21.8 = 257.24 sq cm